OpenStudy (anonymous):

S= 4m-3n -------- 21 solve the ewuation for m

OpenStudy (sriram):

[21*S+3n]/4

jhonyy9 (jhonyy9):

21S =4m-3n 4m=21S+3n m=(21S+3n)/4 = 3(7S+n)/4